# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v2

@@ 3.2 Well-posedness
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{3}{2}-\frac{\gamma}{2}}}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{H^{\frac{\gamma}{2}}},$
@@
Finally, we give a proof of Theorem 1.2
@@
Firstly, we claim that there exist $v$ satisfying equation (3.3) and energy estimate (3.7). In fact, choosing a function $\phi:\mathbb{R}\mapsto\mathbb{R}_{+}$, that is even, smooth, compactly supported, decreasing on $\mathbb{R}_{+}$, and such that $\phi(0)=1$, and then setting
And $X^{\epsilon}$ solves the equation
Without loss of generality, we assume that $\|X^{\epsilon}\|_{C_{T}W^{\alpha}}\leq\|X\|_{C_{T}W^{\alpha}}$ and $\|u^{\epsilon}_{0}\|_{H^{s}}\leq\|u_{0}\|_{H^{s}}$. Let $v^{\epsilon}$ be the solution of equation
and $T^{*}$ be the time given by Theorem 1.1. Then $v^{\epsilon}$ satisfies the condition of Proposition 3.2 on $0,T^{*}]$. Noting that $X\in C_{T}W^{\alpha}$ for $s+1-\frac{\gamma}{2}<\alpha<\frac{\gamma}{2}-\frac{1}{2}$ and $s<\gamma-\frac{3}{2}$ and any $T>0$. We get that $v^{\epsilon}$ is uniformly bounded in $C_{T^{*}}L^{2}$. Then by Banach–Alaoglu Theorem, Fatou’s lemma([BCD11]) and Proposition 3.2, there exists $\psi:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ and a subsequence $v_{\phi(\epsilon)}$ satisfying
$\lim_{\epsilon\to 0}\psi(\epsilon)\to 0,\quad\lim_{\epsilon\to 0}v_{\phi(\epsilon)}\to v\quad in\>L^{2},\quad\|v\|_{C_{T}L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}s\leq C(\|u_{0}\|_{L^{2}},\|X\|_{C_{T}W^{\alpha}}).$
To prove $v$ is the solution of (3.3), we need the strong convergence of $v_{\phi(\epsilon)}$ in $L^{2}$ to make sure that $\partial_{x}(v_{\phi(\epsilon)})^{2}$ can be weak-converge to $\partial_{x}(v^{2})$(it’s already valid for $\partial_{x}(X^{\epsilon})^{2}$ and $\partial_{x}(X^{\epsilon}v^{\epsilon})$). We use the Aubin-Lions-Simon Theorem in [Sim86]. It is suffice to show the following two conditions
$\partial_{t}v\in L_{T}^{p}(H^{-\sigma}),\quad and\quad v\in C_{T}H^{s},$ (3.13)
be hold for some $\sigma,s>0$, $p>1$. Observing that $\partial_{t}v=\Lambda^{\gamma}v+\partial_{x}(v^{2})+2\partial_{x}(vX)+\partial_{x}(X^{2})$. By Lemma 2.1 and Proposition 3.2, we have $\partial_{t}v\in L_{T}^{2}H^{-\sigma}$ for any $\sigma\geq 1$. The second condition of (3.13) is a consequence of Lemma (3.4). Then we prove $v$ is the solution of (3.3) which prove our claim.
To obtain the global solution, we proceed by our induction argument. Fix $\alpha\in(\frac{1}{4},\frac{1}{2})$, $\bar{T}\in[0,\infty)$. Choosing $T^{(0)}\leq\bar{T}\wedge T^{*}$ sufficiently small so that the right hand side of (3.2) is bounded by $\|u_{0}\|_{L^{2}}$($T^{(0)}$ can be only dependent on $\|X\|_{C_{\bar{T}}W^{\alpha}}$ and $\|u_{0}\|_{L^{2}}$). Consider equation (3.3) starting from $v(\frac{T^{(0)}}{2})$. Setting $v^{(1)}=v(\frac{T^{(0)}}{2}+t)$, $X^{(1)}=X(\frac{T^{(0)}}{2}+t)$, we find that $v^{(1)}(t)$ satisfies the equation
$\partial_{t}v^{(1)}(t)-\Lambda^{\gamma}v^{(1)}(t)=-\partial_{x}((v^{(1)}+X^{(1)})^{2}),\quad v^{(1)}(0)=v(\frac{T^{(0)}}{2}),$ (3.14)
then we can get a solution $v^{(1)}$ of equation on $0,T^{(0)}\wedge\bar{T}]$ by Proposition 3.1 since $\|v(\frac{T^{(0)}}{2})\|_{L^{2}}\leq\|u_{0}\|_{L^{2}}$ and for any $t\in[0,T^{*}-\frac{3T^{(0)}}{2}\wedge\bar{T}]$, we have
$\|X^{(1)}\|_{C_{t}W^{\alpha}}=\|X\|_{C_{[\frac{T^{(0)}}{2},\frac{T^{(0)}}{2}+t]}W^{\alpha}}\leq\|X\|_{C_{\bar{T}}W^{\alpha}}.$
Repeating such process, we can get $v^{(n)},X^{(n)}$ satisfy for any $t\in[0,\bar{T}-\frac{n+2}{2}T^{(0)}\wedge\bar{T}]$,
$\|v^{(n)}(\frac{T^{0}}{2})\|_{L^{2}}\leq\|u_{0}\|_{L^{2}},\quad\|X^{(n)}\|_{C_{t}W^{\alpha}}\leq\|X\|_{C_{T^{*}}W^{\alpha}},\quad v^{(n)}(t)=v^{(n-1)}(\frac{T^{0}}{2}+t),$ (3.15)
@@
where $v^{(n)}$ on $0,T^{(0)}]$ is also a solution of (3.3). Setting $v:=v^{(0)}$ on $0,\frac{T^{(0)}}{2}]$, $v^{new}:=v^{(i)}$ on $\frac{iT^{(0)}}{2},\frac{(i+1)T^{(0)}}{2}]$ for $i\leq n$, then $v^{new}$ is a solution in $C([0,\frac{(n+1)T^{(0)}}{2}];L^{2}(\mathbb{T}))\cap L^{2}([0,\frac{(n+1)T^{(0)}}{2}];H^{\frac{\gamma}{2}}(\mathbb{T}))$, and $\|v^{new}(\frac{(n+1)T^{(0)}}{2})\|_{L^{2}}\leq\|v(\frac{T^{(0)}}{2})\|_{L^{2}}$. Choosing $n=[\frac{2\bar{T}}{T^{(0)}}]$, we get $v\in C([0,\bar{T}];L^{2}(\mathbb{T}))\cap L^{2}([0,\bar{T}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Since $\bar{T}$ is arbitrary, we obtain $v$ as a weak global solution of (3.3), since $X$ satisfy (1.5), we prove that $v$ is the weak global solution of (1.1).
Finally, we prove that for $\gamma>\frac{5}{3}$, $v$ can be the strong global solution of (1.1). In fact, it’s suffice to show the uniqueness of $v$. Let $u_{0}^{1}$, $u_{0}^{2}$ belongs to $H^{\frac{1}{2}+\delta}$ for $\delta$ satisfying Proposition 3.1, and $v(t;u_{0})$ be the solution of (3.3) with $v(0)=u_{0}$. Let $v^{1}=v(t;u_{0}^{1})$, $v^{2}=v(t;u_{0}^{2})$ and $\overline{v}:=v(t;u^{1}_{0})-v(t;u^{2}_{0})$, then $\overline{v}$ satisfies the equation
Similar as the proof of Proposition 3.2, we obtain the estimates
By Hölder’s inequality, Parsavel formula and Lemma 2.1, for $\gamma>\frac{5}{3}$ and $\gamma-\frac{3}{2}>s>1-\frac{\gamma}{2}$, we have
$\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))\rangle_{0}\leq$ $\|\overline{v}\|_{H^{\frac{\gamma}{2}}}\|\overline{v}(v^{1}+v^{2})\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $\|\bar{v}\|^{1+\frac{1}{\gamma}+}_{H^{\frac{\gamma}{2}}}\|\bar{v}\|^{1-\frac{1}{\gamma}-}_{L^{2}}\|v^{1}+v^{2}\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $\frac{\nu}{2}\|\bar{v}\|^{2}_{H^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|v^{1}+v^{2}\|^{\frac{2\gamma}{\gamma-1}+}_{H^{s}}$
which can be choose by Lemma 3.4. Similarly,
$\langle\overline{v},2\partial_{x}(\overline{v}X))\rangle_{0}\leq\frac{\nu}{2}\|\bar{v}\|^{2}_{H^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|X\|^{\frac{2\gamma}{\gamma-1}+}_{W^{\alpha}}.$
By Gronwall lemma, we get $\|\overline{v}(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 0$ for any $t\geq 0$, which finish our proof. ∎
@@ 3.3 Application to DP equation
For $\gamma\in(\frac{3}{2},2]$, $\alpha\in(\frac{1}{4},\frac{\gamma}{2}-\frac{1}{2})$, $\delta$ satisfying Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$. Let $v$ be a smooth solution of (3.16) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ and $\tilde{X}(t)\in C_{\bar{T}}W^{\alpha}$, then we have the following energy estimate
@@
$\int_{\mathbb{T}}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x\leq$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|\tilde{X}\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}}+C_{\nu}\|\tilde{X}\|_{W^{\alpha}}^{4}.$ (3.20)
@@
$\|v(t)\|_{L^{2}}^{2}+(8-8\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 16\|u_{0}\|_{L^{2}}^{2}+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$ (3.23)
for $\nu<1$ and $t\in[0,T]$, by Gronwall’s inequality, we finish our proof. ∎
Similar as the proof of Lemma 3.4, we can improve the estimate in higher regularity as follows, which we omit the proof.
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $s<\gamma-\frac{3}{2}$, $\delta$ satisfy Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$, let $v$ be the smooth solution of (3.16) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
@@
$+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v(t^{\prime\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime},$
for some small $\nu<1$ and any $0<s<\gamma-\frac{3}{2}$. Moreover, $v\in C_{T}H^{s}\cap L_{T}^{2}H^{s+\frac{\gamma}{2}}$.
Using the similar argument in the proof of Theorem 1.2, we finish the proof of Remark 1.3.

### v3

@@ 3.2 Well-posedness
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}},$
@@
Now we prove that $v$ given by Proposition 3.2 satisfied the energy estimates.
Lemma 3.5.
Let $u_{0}\in\dot{H}^{\sigma}$ for $\sigma>0$, $\gamma>\frac{5}{3}$, then for $\tilde{X}=X$ satisfying (1.5), there exists a unique weak solution $v$ of equation (3.3) satisfying energy estimate (3.14) and blow up criterion (3.9).
Proof.
Noting that $X$ satisfies (1.5), by Proposition 2.7, we have $X\in C_{T}W^{\alpha}$ for any $\alpha\in(\frac{1}{3},\frac{\gamma}{2}-\frac{1}{2})$. Then for $\gamma>\frac{5}{3}$ and $s\in[0,\gamma-\frac{3}{2})$, we can choose $\alpha$ satisfying
$\alpha\geq\max(\frac{3}{2}-\frac{3\gamma}{4},s+1-\frac{\gamma}{2}).$
Choose a function $\phi:\mathbb{R}\mapsto\mathbb{R}_{+}$ that is even, smooth, compactly supported, decreasing on $\mathbb{R}_{+}$, and such that $\phi(0)=1$, and then set
Let $X^{\epsilon}$ solve the equation
Without loss of generality, we assume that $\|X^{\epsilon}\|_{C_{T}W^{\alpha}}\leq\|X\|_{C_{T}W^{\alpha}}$ and $\|u^{\epsilon}_{0}\|_{\dot{H}^{s}}\leq\|u_{0}\|_{\dot{H}^{s}}$. Let $v^{\epsilon}$ be the solution of equation
Then by Proposition 3.2, there exists $T_{u_{0}}$(Noting that $T_{u_{0}^{\epsilon}}\geq T_{u_{0}}$.) such that $v^{\epsilon}\in L^{4}_{T_{u_{0}}}\dot{H}^{s+\frac{\gamma}{4}}$ for any $s\geq 0$. Noting that $X^{\epsilon}\in C_{T}W^{\infty}$, $u_{0}^{\epsilon}\in\dot{H}^{\infty}$, by Lemma 3.3, we have $v^{\epsilon}\in C_{T}\dot{H}^{s}\cap L_{T}^{2}\dot{H}^{s+\frac{\gamma}{2}}$ for any $s\geq 0$. Choosing $s=0$, then by Banach–Alaoglu Theorem, Fatou’s lemma([BCD11]) and Proposition 3.3, there exists $\psi:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ and a subsequence $v_{\psi(\epsilon)}$ satisfying
$\lim_{\epsilon\to 0}\psi(\epsilon)\to 0,\quad\lim_{\epsilon\to 0}v_{\psi(\epsilon)}\rightharpoonup v\quad in\>L^{2},\quad\|v\|_{C_{T}L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}s\leq C(\|u_{0}\|_{L^{2}},\|X\|_{C_{T}W^{\alpha}}).$
To prove $v$ is the solution of (3.3), we need the strong convergence of $v_{\psi(\epsilon)}$ in $L^{2}$ to make sure that $\partial_{x}(v_{\psi(\epsilon)})^{2}$ can be weak-converge to $\partial_{x}(v^{2})$(it’s already valid for $\partial_{x}(X^{\epsilon})^{2}$ and $\partial_{x}(X^{\epsilon}v^{\epsilon})$). We use the Aubin-Lions-Simon Theorem in [Sim86]. It suffices to show the following two conditions
$\partial_{t}v^{\epsilon}\in L_{T}^{p}(H^{-\sigma}),\quad and\quad v^{\epsilon}\in C_{T}H^{s},$ (3.20)
uniformly hold for some $\sigma,s>0$, $p>1$. Observing that $\partial_{t}v^{\epsilon}=\Lambda^{\gamma}v^{\epsilon}+\partial_{x}(v^{\epsilon})^{2}+2\partial_{x}(v^{\epsilon}X^{\epsilon})+\partial_{x}(X^{\epsilon})^{2}$, by Lemma 2.1 and Proposition 3.3, we have $\partial_{t}v^{\epsilon}$ is uniformly bounded for any $\sigma\geq 1$. The second condition of (3.20) is a consequence of Lemma (3.4) when we choose $0<s\leq\sigma$.
Now we get a solution $v$ in $C_{T}L^{2}\cap L_{T}^{2}\dot{H}^{\frac{\gamma}{2}}$ for some $T>0$ and satisfying the energy estimate. Noting that
$\|v(t)\|_{L_{T}^{4}\dot{H}^{\frac{\gamma}{4}}}\leq\|v(t)\|_{L_{T}^{4}H^{\frac{\gamma}{4}}}\leq\|v(t)\|_{C_{T}L^{2}}^{\frac{1}{2}}\|v(t)\|_{L_{T}^{2}H^{\frac{\gamma}{2}}}^{\frac{1}{2}},$ (3.21)
we have that $v$ is also the solution in $L_{T}^{4}\dot{H}^{\frac{\gamma}{4}}$ with initial data in $L^{2}$. By the uniqueness of Proposition 3.2, we get $v$ also has the blow up criterion. Finally, we prove that the chooise of $v$ is independent by the choosing of mollifier $\phi$. It suffices to show the uniqueness of $v$ in $C_{T}L^{2}\cap L^{2}\dot{H}^{\frac{\gamma}{2}}$. Let $u_{0}$ belong to $H^{\sigma}$ for $\sigma>0$ and $v^{1}(t;u_{0})$ and $v^{2}(t;u_{0})$ be the solutions on $[0,T]$ given above and $\overline{v}:=v(t;u^{1}_{0})-v(t;u^{2}_{0})$, then $\overline{v}$ satisfies the equation
Noting that $v^{1}$ and $v^{2}$ belong to $C_{T}L^{2}\cap L_{T}^{2}\dot{H}^{\frac{\gamma}{2}}$, multiplying $\bar{v}$ both sides and integrating, we obtain the estimate
By Hölder’s inequality, Parsavel formula, Lemma 2.1 and (3.7), for $\gamma>\frac{5}{3}$, we have
$\langle\overline{v},\partial_{x}(\overline{v}(v^{1}+v^{2}))\rangle_{0}\leq$ $\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}(v^{1}+v^{2})\|_{\dot{H}^{\frac{\gamma}{2}-\frac{1}{2}}}$
$\leq$ $\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{4}}}\|(v^{1}+v^{2})\|_{\dot{H}^{\frac{\gamma}{4}}}$
$\leq$ $\|\bar{v}\|^{\frac{3}{2}}_{\dot{H}^{\frac{\gamma}{2}}}\|\bar{v}\|^{\frac{1}{2}}_{L^{2}}\|v^{1}+v^{2}\|_{\dot{H}^{\frac{\gamma}{4}}}$
$\leq$ $\frac{\nu}{2}\|\bar{v}\|^{2}_{\dot{H}^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|v^{1}+v^{2}\|^{4}_{\dot{H}^{\frac{\gamma}{4}}}.$
Similarly, by the same estimate as in (3.11), we have
$\langle\overline{v},2\partial_{x}(\overline{v}X))\rangle_{0}\leq$ $2\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}X\|_{\dot{H}^{1-\frac{\gamma}{2}}}$
$\leq$ $\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{2}}}\|\overline{v}\|_{\dot{H}^{\frac{\gamma}{4}}}\|X\|_{W^{\alpha}}$
$\leq$ $\|\bar{v}\|^{\frac{3}{2}}_{\dot{H}^{\frac{\gamma}{2}}}\|\bar{v}\|^{\frac{1}{2}}_{L^{2}}\|X\|_{W^{\alpha}}$
$\leq$ $\frac{\nu}{2}\|\bar{v}\|^{2}_{\dot{H}^{\frac{\gamma}{2}}}+\|\bar{v}\|^{2}_{L^{2}}\|X\|^{4}_{W^{\alpha}}.$
Noting that $\int_{0}^{t}\|v^{1}\|_{\dot{H}^{\frac{\gamma}{4}}}^{4}dt$ and $\int_{0}^{t}\|v^{2}\|_{\dot{H}^{\frac{\gamma}{4}}}^{4}dt$ for $t\in[0,T]$ are finite, by the Gronwall lemma, we get $\|\overline{v}(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|\overline{v}(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 0$ for any $t\in[0,T]$ until the blow-up time, which finishes our proof.
@@
Noting that $u_{0}\in\dot{H}^{\sigma}$ for $\sigma>0$, let $v$ be the solution satisfying Lemma 3.5. To obtain the global solution, we proceed with our induction argument. Firstly, find a $T_{0}$ satisfying
$CT_{0}^{\frac{1}{4}}\|X\|_{C_{\bar{T}}W^{\alpha}}\leq\frac{1}{4}.$
We prove that for any $u_{0}\in H^{\sigma}$ for some $\sigma>0$, there exists a local solution on $0,T_{0}]$. Indeed, by Proposition 3.2, there exists $v$ on $0,T_{u_{0}}\wedge T_{0})$ satisfying energy estimate (3.14) and the following blow-up criterion that if $T_{u_{0}}\leq T_{0}$, then
$\int_{0}^{T_{u_{0}}}\|v(t)\|^{4}_{\dot{H}^{\frac{\gamma}{4}}}dt=\infty.$
By interpolation (3.21), if $T_{u_{0}}\leq T_{0}$, then we have
$\forall T<T_{u_{0}},\quad\int_{0}^{T}\|v(t)\|^{4}_{\dot{H}^{\frac{\gamma}{4}}}dt\leq C(\|u_{0}\|_{H^{\sigma}},\|X\|_{C_{\bar{T}}W^{\alpha}},T_{0})<\infty.$
By contradiction, we have $T_{u_{0}}>T_{0}$, which shows that $v$ is the solution on $0,T_{0}]$. Consider equation (3.3) with initial data $v(T_{0})$. Setting $v^{(1)}(t)=v(T_{0}+t)$, $X^{(1)}(t)=X(T_{0}+t)$, we find that $v^{(1)}(t)$ satisfies the equation
$\partial_{t}v^{(1)}(t)-\Lambda^{\gamma}v^{(1)}(t)=\partial_{x}((v^{(1)}+X^{(1)})^{2}),\quad v^{(1)}(0)=v({T_{0}}).$ (3.22)
By a similar argument, we can also get a solution $v^{(1)}$ of equation on $0,T_{v(T_{0})}\wedge T_{0}[$ , since
$\|X^{(1)}\|_{C_{t}W^{\alpha}}=\|X\|_{C_{[T_{0},T_{0}+t]}W^{\alpha}}\leq\|X\|_{C_{\bar{T}}W^{\alpha}},\quad for\>t\in[0,\bar{T}-T_{0}].$
Then using the similar argument as above implies that $v^{(1)}(t)$ is the solution on $0,T_{0}]$. Repeating such process, we can get $v^{(n)},X^{(n)}$ satisfying
$\|X^{(n)}\|_{C_{t}W^{\alpha}}\leq\|X\|_{C_{\bar{T}}W^{\alpha}},\quad v^{(n)}(t)=v^{(n-1)}(T_{0}+t),\quad for\>t\in[0,\bar{T}-nT_{0}],$ (3.23)
@@
where $v^{(n)}$ on $0,T_{0}]$ is also a solution of (3.3). Setting $v:=v^{(0)}$ on $0,T_{0}]$, $v^{new}:=v^{(i)}$ on $iT_{0},(i+1)T_{0}]$ for $i\leq n$, $v^{new}$ is a solution in $C([0,(n+1)T_{0}];L^{2}(\mathbb{T}))\cap L^{2}([0,(n+1)T_{0}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Choosing $n=[\frac{\bar{T}}{T_{0}}]$, we get $v\in C([0,\bar{T}];L^{2}(\mathbb{T}))\cap L^{2}([0,\bar{T}];H^{\frac{\gamma}{2}}(\mathbb{T}))$. Since $\bar{T}$ is arbitrary, we obtain $v$ as a weak global solution of (3.3). By the proof of Lemma 3.5, it’s unique. Since $X$ satisfies (1.5), we prove that $v$ is the weak global solution of (1.1). ∎
@@ 3.3 Application to DP equation
For $\gamma\in(\frac{3}{2},2]$, $\alpha\in(1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ and $0\leq T\leq\bar{T}<\infty$. Let $v$ be a smooth solution of (3.24) on $[0,T]$ for $u_{0}\in L^{2}$ and $\tilde{X}(t)\in C_{\bar{T}}W^{\alpha}$, then we have the following energy estimate
@@
$\int_{\mathbb{T}}\partial_{x}(2v\tilde{X}+\tilde{X}^{2})w\mathrm{d}x\leq$ $\frac{\nu}{3}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{L^{2}}^{2}\|\tilde{X}\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}+C_{\nu}\|\tilde{X}\|_{W^{\alpha}}^{4}.$ (3.28)
@@
$\|v(t)\|_{L^{2}}^{2}+(8-8\nu)\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq 16\|u_{0}\|_{L^{2}}^{2}+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{2\gamma}{\gamma-1}}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}$ (3.31)
for $\nu<1$ and $t\in[0,T]$. By Gronwall’s inequality, we finish our proof. ∎
Similar to the proof of Lemma 3.4, we can improve the estimate in higher regularity as follows. The proof is omitted.
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $0<s<\gamma-\frac{3}{2}$, and $0\leq T\leq\bar{T}<\infty$. Let $v$ be the smooth solution of (3.24) on $[0,T]$ for $u_{0}\in H^{s}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
@@
$+\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v(t^{\prime\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$
for some small $\nu<1$. Moreover, $v\in C_{T}H^{s}\cap L_{T}^{2}H^{s+\frac{\gamma}{2}}$.
Using a similar argument as in the proof of Theorem 1.2, we finish the proof of Remark 1.3.